\begin{thebibliography}{10}

\bibitem{Aalst2005}
W.~M. P. V.~D. Aalst, M.~Weske, and D.~Grunbauer.
\newblock {Case Handling: A New Paradigm for Business Process Support}.
\newblock {\em Data and Knowledge Engineering}, 53(2):129--162, 2005.

\bibitem{Bohringer2011}
M.~B\"{o}hringer.
\newblock {Emergent Case Management for Ad-hoc Processes: A Solution Based on
  Microblogging and Activity Streams}.
\newblock In M.~zur Muehlen and J.~Su, editors, {\em Business Process
  Management Workshops}, volume~66 of {\em Lecture Notes in Business
  Information Processing}, pages 384--395. Springer Berlin Heidelberg, 2011.

\bibitem{Clair2009}
L.~C. Clair, C.~Moore, and R.~Vitti.
\newblock {Dynamic Case Management — An Old Idea Catches New Fire}.
\newblock Technical report, Forrester, Cambridge, MA, 2009.

\bibitem{Davenport1994}
T.~Davenport and N.~Nohria.
\newblock {Case Management and the Integration of Labor}.
\newblock {\em MIT Sloan Management Review}, 35(2):11–23, 1994.

\bibitem{deMan2010}
H.~{de Man}, S.~Prasad, and T.~{van Donge}.
\newblock {\em {Mastering the Unpredictable: How Adaptive Case Management Will
  Revolutionize the Way That Knowledge Workers Get Things Done}}, chapter
  {Innovation Management}, pages 211--255.
\newblock Meghan-Kiffer Press, Tampa, Florida, USA, 1st edition, 2010.

\bibitem{DiCiccio2014}
C.~{Di Ciccio}, A.~Marrella, and A.~Russo.
\newblock {Knowledge-Intensive Processes: Characteristics, Requirements and
  Analysis of Contemporary Approaches}.
\newblock {\em Journal on Data Semantics}, Apr. 2014.

\bibitem{Kaan2006}
K.~Kaan, H.~a. Reijers, and P.~V.~D. Molen.
\newblock {Introducing Case Management: Opening Workflow Management’s Black
  Box}.
\newblock In S.~Dustdar, J.~Fiadeiro, and A.~Sheth, editors, {\em BPM 2006,
  LNCS}, volume 4102, pages 358--367, 2006.

\bibitem{Kerremans2008}
M.~Kerremans.
\newblock {Case Management Is a Challenging BPMS Use Case}.
\newblock Technical Report December, Gartner, Stamford, CT, 2008.
\newblock Gartner research number G00162739.

\bibitem{McCauley2010}
D.~McCauley.
\newblock {\em {Mastering the Unpredictable: How Adaptive Case Management Will
  Revolutionize the Way That Knowledge Workers Get Things Done}}, chapter
  {Achieving Agility}, pages 257--275.
\newblock Meghan-Kiffer Press, Tampa, Florida, USA, 1st edition, 2010.

\bibitem{Motahari2013}
H.~R. Motahari-Nezhad and K.~D. Swenson.
\newblock {Adaptive Case Management: Overview and Research Challenges}.
\newblock In {\em Proceedings of the 2013 IEEE 15th Conference on Business
  Informatics}, CBI '13, pages 264--269, Washington, DC, USA, 2013. IEEE
  Computer Society.

\bibitem{Palmer2010}
N.~Palmer.
\newblock {\em {Mastering the Unpredictable: How Adaptive Case Management Will
  Revolutionize the Way That Knowledge Workers Get Things Done}}, chapter
  {Introduction}, pages 1--4.
\newblock Meghan-Kiffer Press, Tampa, Florida, USA, 1st edition, 2010.

\bibitem{Swenson2010}
K.~D. Swenson, editor.
\newblock {\em {Mastering the Unpredictable: How Adaptive Case Management Will
  Revolutionize the Way That Knowledge Workers Get Things Done}}.
\newblock Meghan-Kiffer Press, 1st edition, 2010.

\bibitem{Swenson2010a}
K.~D. Swenson.
\newblock {\em {Mastering the Unpredictable: How Adaptive Case Management Will
  Revolutionize the Way That Knowledge Workers Get Things Done}}, chapter {The
  Nature of Knowledge Work}, pages 5--28.
\newblock Meghan-Kiffer Press, Tampa, Florida, USA, 1st edition, 2010.

\bibitem{Swenson2013}
K.~D. Swenson.
\newblock {State of the Art in Case Management}.
\newblock white paper, Fujitsu, 2013.

\bibitem{Swenson2015}
K.~D. Swenson and J.~Farris.
\newblock {\em {When Thinking Matters in the Workplace}}.
\newblock Purple Hills Books, San Jose, CA, USA, February 2015.

\bibitem{White2009}
M.~White.
\newblock {Case Management: Combining Knowledge With Process}.
\newblock Technical report, BPTrends, July 2009.

\end{thebibliography}

\documentclass[letterpaper]{article}
\pdfoutput=1
\usepackage{hyperref}
\usepackage[english]{babel}
\usepackage[square,sort,comma,numbers]{natbib}
\usepackage{placeins}
\usepackage{tabu}
\usepackage{longtable}
\usepackage{etoolbox}
\usepackage[margin=1in]{geometry}

\hypersetup{colorlinks=true,linkcolor=blue,citecolor=blue,filecolor=blue,urlcolor=blue,
pdftitle={Case Management: a data set of definitions},
pdfauthor={Mike A. Marin, and Matheus Hauder},
pdfsubject={Computer Science, Information systems},
pdfkeywords={Case Handling, Case Management, Case Management Definition}}

\title{Case Management: a data set of definitions}
\author{Mike A. Marin$^{1,2}$, and Matheus Hauder$^{3}$ \\
$^{1}$ IBM Analytics Group\\
Costa Mesa, California, USA\\
$^{2}$ University of South Africa (UNISA)\\
Pretoria, South Africa \\
\texttt{mmarin@acm.org}\\
$^{3}$ Technische Universit\"{a}t M\"{u}nchen (TUM)\\
Boltzmannstr. 3, 85748 Garching bei M\"{u}nchen, Germany\\
\texttt{matheus.hauder@tum.de}}

\begin{document}
\maketitle

\begin{abstract}
Knowledge-intensive processes (KiPs) are becoming increasingly important for organizations with the rise of the knowledge society. Due to their unpredictable and emergent characteristic worklfow management solutions are not suitable to support KiPs. Various case management related approaches have been proposed by researchers and practitioners to support characteristics of KiPs. In this paper we provide a comprehensive list of definitions available on case management, e.g. case handling, adaptive case management, dynamic case management, production case management. For every definition we present the explicit definition, paragraphs that better describe and summarize the case management approach, or extracted sequences that define the term in the referenced publication. All of these definitions are compared against characteristics of KiPs in order to get about understanding of the domain.
\end{abstract}

%\layout{}

\tableofcontents
%\listoffigures
%\listoftables
\pagebreak
\newpage

\section*{Introduction}
\addcontentsline{toc}{section}{Introduction}
We used Google Scholar, IEEExplore, ACM Digital Library, Google Search, and the library of our research institutions using the terms "case management" and "knowledge-intensive process" and trying to identify the first papers that provided definition. Papers in the legal or health care fields were excluded. 
Three types of definitions were extracted, 
\begin{description}
\item[Explicit definitions.] If the paper had an explicit definition that was used.
\item[Paragraphs.] Papers without formal definitions, the paragraph that better described and summarized case management was used as the definition.
\item[Extracted sentences.] Papers without a paragraph that could be used as definition, sentences that defined the term were extracted.
\end{description}
The resulting definitions are presented in the next section.
For each definition, we break it down in what seems to be the main sentences or ideas, and in a set of concepts or components. We also compared the definition with the C1 to C8 KiP characteristics identified by Ciccio \textit{et al.} \cite{DiCiccio2014}.

\section*{Definitions}
\addcontentsline{toc}{section}{Definitions}

\begin{longtabu}{p{16.5cm}}  %\hline
\centering
%%%%%%%%%%%%%%%%%%%%%% BEGIN TABLE %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%\begin{table}
%\centering
% [inline block 0: 24 envs, 27095 chars -> data_tex | \begin{tabular}{|p{1.2cm}|p{5cm}|p{9.23cm}|} \hline 1994 ...]

%\caption{2009, Case Management by White \cite{White2009}}
%\label{table:White2009}
%\end{table}
%%%%%%%%%%%%%%%%%%%%%% END TABLE %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
~ \\
\addcontentsline{toc}{subsection}{2010 Mastering the Unpredictable \cite{Swenson2010}}
\textbf{2010 Mastering the Unpredictable \cite{Swenson2010}} \\
The following five definitions were extracted from Swenson \cite{Swenson2010} \\
~ \\
~ \\
%~ \\ \hline
%~ \\
%%%%%%%%%%%%%%%%%%%%%% BEGIN TABLE %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%\begin{table}
%\centering
% [inline block 1: 32 envs, 32185 chars -> data_tex | \begin{tabular}{|p{1.2cm}|p{5cm}|p{9.23cm}|} \hline 2010 ...]

%\caption{2013, Production Case Management by Swenson \cite{Swenson2013} and \cite{Motahari2013,Swenson2015}}
%\label{table:Swenson2013a}
%\end{table}
%%%%%%%%%%%%%%%%%%%%%% END TABLE %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
~ \\ %\hline
\end{longtabu}

\FloatBarrier
\pagebreak
\newpage
\addcontentsline{toc}{section}{References}
\bibliographystyle{abbrv}
\bibliography{CMMN-KiP}

\end{document}